OpenStudy (anonymous):

determine between which consecutive integers the real zeros of f(x)=2x^2-5x+1 are located

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

there are two real zeros, one at \(\frac{5-\sqrt{17}}{4}\) about .2 which is between 0 and 1 the other is \(\frac{5+\sqrt{17}}{2}\) about 2.28 between 2 and 3